OpenStudy (anonymous):

Find the indefinite integral of: x-4 --------- (x^2 + 2)^2 dx

OpenStudy (anonymous):

I would like a hint please, not the answer.

OpenStudy (anonymous):

take x^2+2 = t That's the hint

OpenStudy (turingtest):

partial fractions is the only hint I can think of ...^that sub wasn't working for me, but maybe I messed up

OpenStudy (turingtest):

also a trig sub could work

OpenStudy (turingtest):

...but that would be ugly

OpenStudy (anonymous):

And write it separately like \[\int\limits_{}^{} ({{x}\over ({x^2+2})^2} - {{4} \over (x^2+2)^2}) dx\]

OpenStudy (anonymous):

Shivam I tried that but how would you do it for the latter term? I realized that I could just use Trig sub and it would work out fine. I get secx's cancelling from the denominator.

OpenStudy (anonymous):

@TuringTest , a trig substitution of x = sqrt(2) tan (theta) can also help

OpenStudy (anonymous):

@QRAwarrior , yes it would get ugly that way. trig substitution seems the best way out

OpenStudy (anonymous):

@QRAwarrior , I confirm I am getting the answer with trig substitution of x = sqrt(2) tan (theta) easily.